OpenStudy (anonymous):

S varies directly as T. If S is 20 when T is 4, then T is when S is 30.

OpenStudy (anonymous):

\[S = k \times T\] S = 20 when T = 4 Find k from here. \[k = \frac{20}{4} = 5\] Now you have k = 5 and S = 30 Find T now: \[T = \frac{S}{k} = \frac{30}{5} = ? \;?\]

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

What??

OpenStudy (anonymous):

what is the anser?

OpenStudy (anonymous):

What you get when you divide 30 by 5??

OpenStudy (anonymous):

If x:6 as 3:9, then x is equal to .

OpenStudy (anonymous):

Firstly tell me answer of first one.. What is the value of T??

OpenStudy (anonymous):

idfknow

OpenStudy (anonymous):

When you divide 30 by 5 you don;t know this?? \[\huge \frac{30}{5}\]

OpenStudy (anonymous):

k thax